OpenStudy (anonymous):

What is the solution for x in the proportion ? show your steps plz 12/22= x/99

OpenStudy (anonymous):

get x by itself multiply both sides by 99

OpenStudy (anonymous):

12/22= x/99 6/11= x/99 x=6 *99 /11 x= 6*9 x=54

OpenStudy (anonymous):

same thing that I just did

OpenStudy (amistre64):

my left side aint refreshing.... who broke what? :)